(3.1)

(a) Find a and b such that −3ai − (−1 − i)b = 3a − 2bi.

(b) Let z1 = 12 + 5i and z2 = (3 − 2i)(2 + λi). Find λ without resorting to division such that z2 = z1


(3.2) Let z = −2 + 3i and z' = 5 − 4i. Determine the complex numbers

(a) z2 − zz'

(b) 1/2(z + z)2

(c) 1/2[z − z] + [(−1 − z')]2.


1
Expert's answer
2021-07-07T14:56:54-0400

(3.1)

"(a)\\
-3ai-(-1-i)b=3a-2bi\\
-3ai+b+bi=3a-2bi\\
(-3a+b)i+b=3a-2bi\\
\therefore\\
-3a+b=-2b\\
b=3a\\
-3a+3a=-2b\\
-2b=0\\
b=0\\
a=0\\
(b)\\
z1 = 12 + 5i \\
z2 = (3 − 2i)(2 + λi)=6+3\lambda{i}-4i+2\lambda\\
z_2=z_1\\
6+3\lambda{i}-4i+2\lambda=12+5i\\
6+2\lambda=12\\
3\lambda{}-4=5\\"

For each last two equation

"\lambda=3"

(3.2)

(a)

"z = −2 + 3i\\
 z' = 5 − 4i\\
z^2=(-2)^2+2(-2)(3i)+(3i)^2=-5-12i\\
zz'=(-2+3i)(5-4i)=-10+8i+15i+12\\
zz'=2+23i\\
z^2-zz'=-7-35i\\"

(b)

"\frac{1}{2(z+z')^2}=\frac{1}{2(-2+3i+5-4i)^2}=\frac{1}{2(3-i)^2}=\frac{1}{2(8-6i)}=\\\frac{1}{16-12i}\\"

Rationalizing

"\frac{16+12i}{(16+12i)(16-12i)}\\
\frac{16+12i}{(16+12i)(16-12i)}\\
\frac{16+12i}{(16^2-(12i)^2)}\\
\frac{16+12i}{(400)}\\
\frac{16+12i}{400}\\
\frac{16}{400}+\frac{12i}{400}\\
\frac{1}{25}+\frac{3i}{100}"

(c)

"\frac{1}{2[z-z']}+[(-1-z')]^2\\
\frac{1}{2[-2+3i-(5-4i)]}+[(-1-(5-4i)]^2\\

\frac{1}{2[-7+7i]}+[-6+4i]^2\\
\frac{1}{-14+14i}+36-48i-16\\
\frac{-14-14i}{(-14+14i)(-14-14i)}+20-48i\\
\frac{-14-14i}{(196+196)}+20-48i\\
\frac{-14-14i}{392}+20-48i\\
\frac{-14}{392}-\frac{14i}{392}+20-48i\\
\frac{-1}{28}-\frac{i}{28}+20-48i\\
\frac{559}{28}-\frac{1345i}{28}"